OpenStudy (anonymous):

Use the quadratic formula to find the roots of the equation. Round to tenths place if necessary. 11x2 - x - 3 = 0 A. {-1, 1.2} B. {-11, 13.2} C. {-0.5, 0.6} D. no real solution

OpenStudy (anonymous):

\[\frac{1\pm\sqrt{133}}{22}\]

OpenStudy (anonymous):

c is closest

OpenStudy (anonymous):

Okay but its not positive? thats weird.. :P

OpenStudy (anonymous):

it is just not exact. it is right to the nearest tenth.

OpenStudy (gw2011):

Quadratic formula: x=[-b+/- square root (b^2-4ac)]/2a a=11 b=-1 c=-3 Substituting these into the formula, you get the answer as: x=[1+/- square root (133)]/22 square root 133=11.5 approximately (1+11.5)/22 and (1-11.5)/22 =12.5/22 and -10.5/22 =0.56 and -0.5 So, the closest answer is (c)=-0.5, 0.6